OpenStudy (anonymous):

Let f(x)=log2(5x)-3. Find f^-1(6).

OpenStudy (anonymous):

x = log2(5y) -3 6 = log 2(5y) -3 9 = log 2(5y)

OpenStudy (anonymous):

so the final answer is 9=log2(5y)?

OpenStudy (anonymous):

No there are more algebra on simplifying it

OpenStudy (anonymous):

so i shoudnt leave it as is?

OpenStudy (anonymous):

no you need to isolate y

OpenStudy (anonymous):

9/(log2(5)) = y

OpenStudy (anonymous):

ok thanks for all the help. i really appreciate it.